OpenStudy (aj_harleyy):

What is an equation that when solved, x= -3

OpenStudy (michele_laino):

the simplest equation is: \(x+3=0 \)

OpenStudy (aj_harleyy):

i need a harder equation that that to equal -3

OpenStudy (michele_laino):

do you mean that x=-3 has to be a solution?

OpenStudy (aj_harleyy):

yes

OpenStudy (michele_laino):

another equation, can be this: \[\log \left( {{x^2} - 8} \right) = 0\]

OpenStudy (michele_laino):

where logarithm is taken in any base

OpenStudy (aj_harleyy):

how would i solve that? I don't know how to do log i just started algebra 1

OpenStudy (alexandervonhumboldt2):

ok here it is: cuberoot(-(x^2+45)/2) here

OpenStudy (michele_laino):

sorry! I write another equation

OpenStudy (aj_harleyy):

okay michele

OpenStudy (michele_laino):

another eqaution is: \[\frac{{x + 3}}{{x - 5}} = 0\]

OpenStudy (aj_harleyy):

thamk you very much!

OpenStudy (michele_laino):

:)